OpenStudy (anonymous):

What is the midpoint of a line segment with endpoints (0,-6) and (-8,0)?

OpenStudy (anonymous):

A. (4, 3) B. (4, -3) C. (-4, -3) D. (-4, 3)

OpenStudy (anonymous):

Mid-Point of points (x1,y1) and (x2,y2) is ( (x1+x2)/2 ,(y1+y2)/2 )

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

which one is x and which one is y?

OpenStudy (anonymous):

or does it matter?

OpenStudy (anonymous):

x1 =0 , x2=-8 , y1 =-6 and y2=0 Now, use Mid-Point of points (x1,y1) and (x2,y2) is ( (x1+x2)/2 ,(y1+y2)/2 )

OpenStudy (anonymous):

(-8)/2 = -4 (-6)/2 = -3

OpenStudy (anonymous):

so it'd be (-4,-3) ?

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (anonymous):

Thank you! And instead of posting another question i'll post it here... cause you help me a lot! this ones similar but kind of backwards... 9. The midpoint of a line segment is (-1,-3). One endpoint is (7,3). What is the other endpoint? how would i do that one?

OpenStudy (anonymous):

here, let the coordinates of other end point be (x1 ,y1)

OpenStudy (anonymous):

So, (x1+7)/2 =-1 AND (y1+3)/2 =-3

OpenStudy (anonymous):

Now, can u find x1 and y1

OpenStudy (anonymous):

x1= -9?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y1 = -9?

OpenStudy (anonymous):

GREAT

OpenStudy (anonymous):

Thanks for all your help!

OpenStudy (anonymous):

welcome